OpenStudy (anonymous):

log[log(x+4)-4]=0 then x=-4 how

OpenStudy (anonymous):

if \[\log(\log(x+4)-4)=0\] then it must be the case that \[\log(x+4)-4=1\]

OpenStudy (anonymous):

that makes \[\log(x+4)=5\] and so \[x+4=10^5\] \(x=-4\) is NOT the answer for sure you cannot take the log of zero